OpenStudy (anonymous):

A particles position as a function of time is given by u = cos 2t i+sin 2t j ( The u, i and j should have tilde's underneath as they are vectors) (a) find the equation of the path

OpenStudy (amistre64):

x = cos(2t) y = sin(2t) z = 0

OpenStudy (anonymous):

thats not the answer i sorta know how to do it but the 2t throws me off abit

OpenStudy (amistre64):

thats the equation of a line in R^3 :) well, except for an initial point

OpenStudy (amistre64):

the position path just produces a unit circle

OpenStudy (anonymous):

yes i understand that but normally you would eliminiate the sin and cos by squaring x and y then adding them

OpenStudy (anonymous):

im not sure what happens to the 2t

OpenStudy (amistre64):

2t just modifies the speed of rotation; nothing more

OpenStudy (amistre64):

its going twice as fast around the unit circle

OpenStudy (anonymous):

i see so the working is like this; x = cos 2t y=sin2t x^2=cos^2t y^2=sin^2t x^2+y^2=1 i ignore the 2t correct

OpenStudy (amistre64):

or simply redefine it; 2t = u

OpenStudy (amistre64):

it has no importance on position; other than it a given t it is at this location; but the path remains unchaged by it

OpenStudy (amistre64):

tell me; does a person walking along the sidewalk get to the same place as the person jogging on the sidewalk?

OpenStudy (anonymous):

yes thats correct. thanks so much you actually made me understand the question.

OpenStudy (amistre64):

youre welcome :)